OpenStudy (anonymous):

pleasee help stuck on this problem :/ ! Chocolate Box Company is going to make open-topped boxes out of 8 × 11-inch rectangles of cardboard by cutting squares out of the corners and folding up the sides. What is the largest volume box it can make this way? (Round your answer to the nearest tenth.)

OpenStudy (anonymous):

so far I have gotten 88-12x+12x^2=0 but I dont know what to do next!

OpenStudy (anonymous):

h=x l=8-2 v=(8-2x)(11-2x)x v=(8-2x)(11x-2x^2) v=88x-6x^2+4x^3 v'=88-12x+12x^2 Those are my previous steps to get there

OpenStudy (anonymous):

You stumped me, sorry.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

its okay, thanks for taking your time to look over it!

OpenStudy (anonymous):

8 × 11 square inches v=l*b*h mae l b and h equal l=b=h THat is one hint

OpenStudy (anonymous):

To find volume don't you need L * W * H?

OpenStudy (anonymous):

yes whoops I forgot to write v=l*w*h

OpenStudy (anonymous):

length base and height but same thing tho

OpenStudy (anonymous):

timo86m sorry I didnt understand that, is my work so far correct because I have done some of the problem as much as I could,just confused with what to do next

OpenStudy (anonymous):

I am not sure what you are going for. BUt A hint is that making a box with all sides the same size is the biggest box yo ucan make.

OpenStudy (anonymous):

yeah the problem is looking for the largest volume the box can make

OpenStudy (fellowroot):

I solved a very similar problem earlier. Here is my solution.